OpenStudy (anonymous):

Solve x/-4 - (-8) = 12. please help me

OpenStudy (shamil98):

\[\frac{ x }{ -4 } + 8 = 12\] subtract 8 from both sides. \[\frac{ x }{ -4 } = 4\] multiply both sides by -4 x = -16

OpenStudy (anonymous):

x=-16 thats the answer i got

OpenStudy (anonymous):

I agree

OpenStudy (anonymous):

It is x = -16! Solve it... -16/-4 = 4. 4 - (-8) = 12. So it is -16